8. Nonconstant growth stock Aa Aa As companies evolve, certain factors can drive sudden growth. This may lead to a period of nonconstant, or variable, growth. This would cause the expected growth rate to increase or decrease, thereby affecting the valuation model. For companies in such situations, you would refer to the variable, or nonconstant, growth model for the valuation of the companys stock. Consider the case of Portman Industries: Portman Industries just paid a dividend of $2.16 per share. The company expects the coming year to be very profitable, and its dividend is expected to grow by 16.00% over the next year. After the next year, though, Portmans dividend is expected to grow at a constant rate of 3.20% per year. The risk-free rate (Rr) is 4.00%, the market risk premium (RPM) is 4.80%, and Portmans beta is 2.00. Term Value Dividends one year from now (D1) Horizon value (P1) Intrinsic value of Portmans stock Assuming that the market is in equilibrium, use the information just given to complete the table What is the expected dividend yield for Portmans stock today? o 8.32% О 10.40% 0 10.08% О 11.45%
Answer #1

Dividends one year from now (D1) = D*(1+g1) = 2.16*(1+16%) = 2.51

D2 = D1*(1+g1)*(1+g2) = 2.16*(1+16%)*(1+3.2%) = 2.5858 or 2.59

Required Rate = Risk Free rate + Beta *( Market return - Risk Free Rate) = 4% + 2*(4.8%) = 13.6%

Horizon Value value = D2/(Required Rate - Growth) = 2.5858/(13.6%-3.2%) = 24.8633 or 24.86

Intrinsic Value of Share = D1/(1+r) + Horizon Value/(1+r)2 = 2.5858/(1+13.6%) + 24.8633/(1+13.6%) = 24.16

Dividend Yield = 2.51/24.16 = 10.40%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well